# 4Diffusion: Multi-view Video Diffusion Model for 4D Generation

### v1

@@ A.1 More Implementation Details
Datasets. We utilize Objaverse [11] to train our multi-view diffusion model, as described in Sec. 3.2. Objaverse dataset comprises a vast collection of 3D shapes with descriptive captions, tags, and animations. We manually filter out animated 3D shapes that contain static objects, out-of-scene movement, rotated cameras, or meaningless objects, resulting in 926 high-quality 3D animated models, as depicted in Fig. 7. We apply Blender to render 32 videos with azimuth angles uniformly ranging from and an elevation angle of for each animated 3D model. In our experiments, we use the dataset released by Consistent4D [19] and text-image pairs from ImageDream [52] project page. Specifically, for text-image pairs, we leverage Stable Video Diffusion V1.1 [5] to produce monocular videos for 4D generation.
@@ A.2 More Results
4D Generation. In Fig. 8, we showcase additional results of our 4D generation results. To gain a more intuitive understanding, we encourage readers to view the supplementary videos.

### v2

@@ 5 Conclusion
Acknowlegements
The work is supported by the National Key R&D Program of China (No. 2022ZD0160102), the National Natural Science Foundation of China under Grant No. 62102150, and the Science and Technology Commission of Shanghai Municipality under Grant No. 23QD1400800.
@@ A.1 More Implementation Details
Datasets. We utilize Objaverse dataset [11] to train our multi-view diffusion model, as described in Sec. 3.2. Objaverse dataset comprises a vast collection of 3D shapes with descriptive captions, tags, and animations. We manually filter out animated 3D shapes that contain static objects, out-of-scene movement, rotated cameras, or meaningless objects, resulting in 926 high-quality 3D animated models, as depicted in Fig. 10. We apply Blender to render 32 videos with azimuth angles uniformly ranging from and an elevation angle of for each animated 3D model. In our experiments, we use the dataset released by Consistent4D [20], test cases from Objaverse, and text-image pairs from ImageDream [52] project page. Specifically, for text-image pairs, we leverage Stable Video Diffusion V1.1 [5] to produce monocular videos for 4D generation.
@@ A.2 More Results
4D Generation. In Fig. 11, we showcase additional results of our 4D generation results. To gain a more intuitive understanding, we encourage readers to view the supplementary videos. Moreover, we show more comparisons on 4D generation as shown in Fig. 12, Consistent4D and DreamGaussian4D encounter the multi-face problem while our method generates spatial-temporally consistent contents.